A 50.00 mL aliquot solution containing 0.4500 g of magnesium sulfate in 0.500 L required 37.60 mL of EDTA solution for titration. How many milligrams of Calcium carbonate will react with 1.00 mL of the same EDTA solution?


1
Expert's answer
2021-02-15T03:18:47-0500

Solution:

EDTA = H2Y2-

magnesium sulfate: MgSO4

The molar mass of MgSO4 is 120.37 g mol-1

Moles of MgSO4 = Mass of MgSO4/Molar mass of MgSO4 = 0.4500 g / 120.37 g mol-1 = 0.0037385 mol

n(MgSO4) = 0.0037385 mol


MgSO4 → Mg2+ + SO42-

According to the equation above: n(MgSO4) = n(Mg2+)

n(Mg2+) = n(MgSO4) = 0.0037385 mol (in 0.5 L of solution)

n(Mg2+) = 0.0037385 mol × (0.050 L / 0.500L) = 0.00037385 (in 50.00 mL of aliquot solution)


Mg2+ + H2Y2- → MgY2- + 2H+

According to the equation above: n(Mg2+) = n(H2Y2-)

n(H2Y2-) = n(Mg2+) = 0.00037385 mol

Moles of H2Y2- = Molarity of H2Y2- × Volume of H2Y2-

Molarity of H2Y2- = Moles of H2Y2- / Volume of H2Y2- = 0.00037385 mol / 0.03760 L = 0.009943 M

Molarity of H2Y2- = 0.009943 M


calcium carbonate: CaCO3

The molar mass of CaCO3 is 100.09 g mol-1

CaCO3 → Ca2+ + CO32-

Ca2+ + H2Y2- → CaY2- + 2H+

According to the equations above: n(H2Y2-) = n(Ca2+) = n(CaCO3)

n(H2Y2-) = n(CaCO3)

Molarity of H2Y2- × Volume of H2Y2- = Mass of CaCO3 / Molar mass of CaCO3

0.009943 M × 0.0010 L = Mass of CaCO3 / 100.09 g mol-1

Mass of CaCO3 = 0.009943 M × 0.0010 L × 100.09 g mol-1 = 0.0009952 g = 0.9952 mg

Mass of CaCO3 = 0.9952 mg CaCO3


Answer: 0.9952 milligrams of CaCO3 will react with 1.00 mL of this EDTA solution.